OpenStudy (anonymous):

complete the chart...fractional notation?... decimal notation 0.4 ....percent notation?

OpenStudy (anonymous):

4/10=2/5

OpenStudy (anonymous):

so 2/5 is percent notation

OpenStudy (anonymous):

oh, percent notation, 40%

Directrix (directrix):

.4 = 4/10 4/10 x 10/10 = 40 / 100 = 40%